OpenStudy (anonymous):

Two trains leave the station at the same time, one heading west and the other east. The westbound train travels at 85 miles per hour. The eastbound train travels at 95 miles per hour. How long will it take for the two trains to be 216 miles apart?

OpenStudy (madhu.mukherjee.946):

The two trains are moving away from each other. train 1 speed = 85 mph train 2 speed = 95 mph.

OpenStudy (madhu.mukherjee.946):

combined speed = 190 mph

OpenStudy (madhu.mukherjee.946):

distance = 216 miles t=d/r t=216/190 t = 1.136 hours After 1.136 hours they will be 216 miles apart

OpenStudy (madhu.mukherjee.946):

we used the formula distance=time *speed

OpenStudy (madhu.mukherjee.946):

@mm41 do you understand?